OpenStudy (anonymous):

I need help to solve Multi-step equations variables on each side. Here's some problems... -45m+68=84m-61 4-2x=-5x-1 0.3x-15=0.2x-5 3/4w-5=1/4w-1 3b+2=6(3-b)

OpenStudy (phi):

How far did you get? This is something you should learn now, because lots of math requires you to know how to do it. If you lose the bubble now, it will be very unpleasant in later math classes.

OpenStudy (anonymous):

Well I got to the part on the P.E.M.D.A.S but after that I got confused. Like on the first problem -45m+68=84m-61 I would have to subtract and divide, I think.

OpenStudy (phi):

in -45m +68 = 84m - 61 I see a -45 m on the left side. I would add +45 m to both sides. that means write +45m on both sides -45m + 45m + 68 = 84m +45m - 61 on the left side you have 45 m's take away 45 m's how many m's are left ?

OpenStudy (anonymous):

0?

OpenStudy (phi):

yes so the left side is 0 + 68 = 84m + 45m - 61 of course, 0+68 is just 68, so you have 68= 84m + 45m - 61 on the right side you have 84 m's plus 45 m's how many m's do you have ?

OpenStudy (anonymous):

129 m's

OpenStudy (phi):

and you should write out the entire equation (good habit to get into) 68 = 129 m - 61 now add +61 to both sides. that means write +61 on both sides can you do that ?

OpenStudy (phi):

68+61 = 129 m - 61+61

OpenStudy (phi):

can you simplify the left side ?

OpenStudy (anonymous):

68+61=129m-0?

OpenStudy (phi):

yes, on the right side. now the left side

OpenStudy (anonymous):

129=129m-0?

OpenStudy (phi):

yes. and adding 0 or subtracting 0 does not "do anything" so you have 129= 129m now divide both sides by 129

OpenStudy (phi):

like this \[ \frac{129}{129}= \frac{129}{129} m \]

OpenStudy (phi):

what is 129 divided by itself ? what is anything divided by itself ?

OpenStudy (anonymous):

1

OpenStudy (phi):

so simplify the equation \[ \frac{129}{129}= \frac{129}{129} m \]

OpenStudy (anonymous):

1=1m

OpenStudy (phi):

yes, and usually we don't write 1m that means 1 times m and 1 times anything is the anything. in other words 1*m = m so we just write 1= m or more usually m=1

OpenStudy (anonymous):

Ohhh okay, thank you

OpenStudy (phi):

can you try the next problem 4-2x=-5x-1 I usually look at the terms with the "x" I would add +5x to both sides. can you do that ?

OpenStudy (anonymous):

4-2x=5x+5x-1 Like that?

OpenStudy (phi):

first, you have to do it both sides (to keep the sides equal) also, the right side should be -5x+5x-1 (you left out a minus sign)

OpenStudy (phi):

4-2x=-5x-1 becomes 4-2x+5x = -5x+5x -1

OpenStudy (anonymous):

Then what do I do. Okay well I have to goo, bye,

OpenStudy (phi):

simplify both sides on the left you have 5 x's take away 2 x's on the right side you have 5 x's take away 5 x's

OpenStudy (anonymous):

4-3=0-1? I think I'm kindaa getting it